OpenStudy (anonymous):

Use the shortcut rules to mentally calculate the derivative of the given function. f(x) = 8x4

OpenStudy (anonymous):

8x^4

OpenStudy (turingtest):

4(8)x^(4-1)=32x^3

OpenStudy (anonymous):

Oh thank you!

OpenStudy (anonymous):

Use the power rule \[f(x) = x^n\] then \[f \prime (x) = nx^{n-1}\]

OpenStudy (anonymous):

And you end up with what Turing said.